OpenStudy (anonymous):

Let f(x)=x^2+x+13. What is the value of x for which the tangent line to the graph of y=f(x) is parallel to the x-axis?

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (xishem):

\[f(x)=x^2+x+13\]\[f'(x)=2x+1\] Any line is parallel to the x-axis when its slope is 0, so find x for when the tangent line's slope is equal to 0: 0 = 2x+1 -1=2x x = -1/2

OpenStudy (anonymous):

y = -2

OpenStudy (anonymous):

@above? how is the line vertical and tangent to that graph at the same time?

OpenStudy (anonymous):

oh i see, maybe, sorry

OpenStudy (xishem):

The line is horizontal, because the x-axis is horizontal. Two horizontal lines are parallel.

OpenStudy (anonymous):

yes i didnt see that it said x-value, i put the actual line that was tangent and parallel

OpenStudy (anonymous):

and i messed that part up too!

OpenStudy (anonymous):

the equation for the tangent line would be y = 51/4

OpenStudy (xishem):

The equation for the tangent line would be: \[y=\frac{-1}{2}x+13\]

OpenStudy (anonymous):

no, plug the axis of symmetry (-1/2) in for x to get 12 3/4

OpenStudy (anonymous):

and yours isnt even horizontal

OpenStudy (xishem):

I apologize. The equation for the tangent line would be: \[y=13\] \[(\frac{-1}{2},\frac{51}{4}) \]is actually the vertex point for the original equation

OpenStudy (anonymous):

y = 13? That would touch it at two points though. You got the vertex, now you just need a horizontal line with the y-coordinate of the vertex so that it touches the vertex.